A coil of wire has a resistance of 10 Ω at 0°C. If the temperature coefficient of resistance for the wire is 0.004/°C, its resistance at 100°C is


1
Expert's answer
2021-08-23T12:40:32-0400

"R_t=R_0(1+\alpha t)=10\cdot(1+0.004\cdot100)=14\ (\Omega)" . Answer